Q(1) Determine Whether Each Of The Following Systems Is Linear: (A) 3x – 4y + 2yz = 8 (b) ex + 3y = 1(C) 2x-3y+kz=4

Q(2)  Write V As A Linear Combination Of U1, U2, And U3 V=(1,4,6) U1=(1,1,2) U2=(2,3,5) U3=(3,5,8)

1
Expert's answer
2021-12-14T18:33:00-0500

1.

Linear systems are systems of equations in which the variables are never multiplied with each other but only with constants and then summed up


a)

not linear: equation has term with variables y, z are multiplied


b)

linear


c)

linear


2.

"v=au_1+bu_2+cu_3"


then:

"a+2b+3c=1"

"a+3b+5c=4"

"2a+5b+8c=6"


if we subtract sum of 1st and 2nd equations from third equation, we get:

"0=1"

so, there are no solutions